OpenStudy (anonymous):

Help? right answer will get a fan and a medal.... If y varies inversely with x and y = 8 when x = 5, then what is y when x = 2?

OpenStudy (zale101):

this is the constant variation y=kx so what you have to do is replace the y with 8 and x with 5 and solve for k 8 = 5k now isolate k can you do that?

OpenStudy (nirmalnema):

x1y1=x2y2 so8*5=2*y y=20

OpenStudy (nirmalnema):

y is inversly proportional. to x